OpenStudy (anonymous):

why does 3600÷6 have the same quotient as 1,800 ÷ 3 ?

OpenStudy (zzr0ck3r):

because 3600*3 = 6* 1800

OpenStudy (anonymous):

\[\frac{ 3600 }{ 6}=\frac{ 1800*2 }{ 3*2 }\]

OpenStudy (anonymous):

Because 1800 is half of 3600 and 3 is half of 6 therefore you get the same answer